# Model-Agnostic Bounds for Augmented Inverse Probability Weighted Estimators’ Wald-Confidence Interval Coverage in Randomized Controlled Trials

### v1

@@ 3.2 Non-cross-fit procedure
    Under Part 1 of Condition 12, the left-hand side of (2) is
    ${\mathrm{O}}\left(\Delta_{{\mathrm{VC}}}\sqrt{\log\{({\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2})^{-1}\}}+n^{-1/2}\log\{({\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2})^{-1}\}+\left\{(\Delta_{{\mathrm{VC}}}+n^{-1/2}\right\}\log n\right),$
    where $\Delta_{{\mathrm{VC}}}$ is $({\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2})^{1/3}$ and can be replaced by $\sqrt{{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}\log n}$ additionally under Conditions 10 and S2 if ${\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}={\mathrm{o}}(1/\log n)$.
@@
    Under Part 2 of Condition 12, the left-hand side of (2) is
    ${\mathrm{O}}\left(\Delta_{{\mathrm{VC-hull}},1}+\Delta_{{\mathrm{VC-hull}},2}+\left\{\Delta_{{\mathrm{VC-hull}},1}^{\nu}+n^{-1/2}\right\}\log n\right),$
    where $\Delta_{{\mathrm{VC-hull}},1}=({\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2})^{1/(1+2\nu)}+n^{-1/(4\nu-2)}$ and $\Delta_{{\mathrm{VC-hull}},2}=0$. Additionally under Conditions 10 and S2, if ${\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}={\mathrm{o}}(1/\log n)$, this rate holds with $\Delta_{{\mathrm{VC-hull}},1}$ replaced by $({\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}\log n)^{1/(2\nu)}$ and $\Delta_{{\mathrm{VC-hull}},2}$ replaced by $n^{-1/2}({\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}\log n)^{-(\nu-1)/\nu}$.
@@ 3.3 Interpretation and comparison of asymptotic rates
For simplicity, particularly when comparing cross-fit and non-cross-fit procedures, I treat $K$ as an absolute constant that does not vary with $n$ for now. Because of the convergence rate of $|\sigma_{\dagger,a}-\sigma_{\#,a}|$ shown in Theorems 2 and 4, the forms in (1) and (2) are somewhat deceiving in the sense that the seemingly first order term $\phi(z_{\alpha})z_{\alpha}(\sigma_{\dagger,a}-\sigma_{\#,a})/\sigma_{\#,a}$ on the left-hand side has a faster convergence rate than the right-hand side, regardless of whether subgaussian conditions hold and whether $Q_{\#,a}$ is chosen to yield a faster rate of $|\sigma_{\dagger,a}-\sigma_{\#,a}|$. In other words, the rate of the right-hand side of (1) and (2) is an upper bound on the convergence rate of the Wald-CI coverage to its nominal coverage $1-\alpha$.
@@
For the cross-fit procedure, the ${\mathrm{O}}(\sqrt{\log n/n})$ term in the rate in Theorem 1 is the same as the rate in the Berry-Esseen Theorem except for a log factor. The contribution from estimating the nuisance outcome model can converge slower and strongly depends on whether subgaussian conditions 7 and S1 hold: If $\hat{Q}_{k,a}$ is estimated from a (potentially misspecified) parametric model, so that ${\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}={\mathrm{O}}(n^{-1})$ where $Q_{\#,a}$ is the mean of $\hat{Q}_{k,a}$, and subgaussian conditions hold, then this contribution ${\mathrm{O}}(\sqrt{{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}}\log n)$ is ${\mathrm{O}}(n^{-1/2}\log n)$, same as the Berry-Esseen theorem except for a log factor. However, if $\hat{Q}_{k,a}$ is estimated more flexibly so that ${\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}$ has a slower rate, or subgaussian conditions do not hold, then the contribution from nuisance estimation can have a slower rate.
@@
For the non-cross-fit procedure, the rate additionally strongly depends on the complexity of the function class $\mathcal{F}$} containing $\tilde{Q}_{a}$, as measured by its covering number. This is because, compared to the cross-fit procedure, Donsker conditions are additionally required to bound empirical process terms, whose rate depends on the function class complexity. The bounds I adopted for empirical process terms might have room for improvement. Among the empirical process terms in my derivation, the term that impacts the rate most is $(P_{n}-P_{*})\{{\mathcal{T}}_{a}(\tilde{Q}_{a})-{\mathcal{T}}_{a}(Q_{\#,a})\}$ concerning the deviation of $\tilde{\psi}_{a}-\psi_{*,a}$ from the sample mean $(P_{n}-P_{*})D_{a}(Q_{\#,a},\psi_{*,a})$. However, as I show in the simulation in Section 5, this term might not severely impact the normal approximation, and my bound on this term might be conservative.
@@
By Theorem 3 and Corollary 1, the rates for VC classes (e.g., many parametric models) are similar to the rates for singleton classes, the simplest possible classes, which achieve the lower bound of the right-hand side of (2). [^1] For singleton classes, the Berry-Esseen Theorem directly applies and yields a tighter bound without log factors, even without subgaussian conditions. These rates are also similar to those for the cross-fit procedure, suggesting that cross-fit and non-cross-fit procedures have similar CI coverage if $\mathcal{F}$} is not rich.
@@
In contrast, the rates for VC-hull classes, which are much richer than VC classes, can be much slower. If the VC index $\nu$ of the associated VC-class is large, namely $\mathcal{F}$} is rich, then the rate would be slow, and the subgaussian conditions do not improve the rate as much as in the cross-fit procedure or the non-cross-fit procedure with VC classes. As an example, suppose that $\mathcal{F}$} is a uniformly bounded subset of the convex hull of $\{x\mapsto c{\mathbbm{1}}(x_{s}\leq u_{s}):u\in{\mathbb{R}}^{d},c\in{\mathbb{R}},s\subseteq\{1,\ldots,d\}\}$, where, for any vector $v\in{\mathbb{R}}^{d}$, $v_{s}$ denotes the entries of a vector $v$ with indices in the index set $s$, and $\leq$ is entrywise. Then any function in $\mathcal{F}$} can be well approximated by a step function, and $\nu=d+2$ (Example 2.6.1 in van der Vaart and Wellner,, 1996). Even for moderately rich covariates such as $d=3$, the bound on the convergence rate would be quite slow, for example, $({\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2})^{1/10}$ or slower (up to log factors). Such rich Donsker classes are common in nonparametric regression. For example, highly adaptive lasso estimator lies in the above example class $\mathcal{F}$} (Benkeser and Van Der Laan,, 2016). Theorem 3 and Corollary 1 are mainly for illustration rather than for proving accurate rates for specific regression algorithms, because tighter bounds might be derived for a specific algorithm to estimate $Q_{*,a}$ with additional knowledge on this algorithm. For example, more accurate characterizations of the complexity for highly adaptive lasso have been obtained (van der Laan,, 2023).
For the cross-fit procedure, the choice of the number of folds $K$ in cross-fitting under exchangeable sample-splitting (Condition 6) is an interesting question. If the impact of $K$ on ${\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}$ is ignored, Theorem 1 suggests choosing a small $K$ because the bound increases with $K$. However, a small $K$ would generally lead to a larger ${\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}$ because of fewer training data. Without additional knowledge or assumptions on how ${\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}$ varies with training sample size n(K1)/Kn(K-1)/K, it appears impossible to obtain concrete theoretical guidance on the choice of $K$. Suppose that ${\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}={\mathrm{O}}(\{n(K-1)/K\}^{-r})$ for some constant $r>0$. For example, if $\hat{Q}_{k,a}$ is estimated in a parametric model, $r$ can be anticipated to be one; if the usual ${\mathrm{o}}_{p}(n^{-1/4})$-rate requirement on $\hat{Q}_{k,a}$ is achieved, $r$ can be anticipated to be larger than 1/2. The resulting rate in Theorem 1 is $n^{-r/3}K^{(r+2)/3}(K-1)^{-r/3}$, or $n^{-r/2}K^{(r+1)/2}(K-1)^{-r/2}$ (up to log factors) under subgaussian conditions 7 and S1. Figure 1 illustrates the multiplicative factor involving $K$ in the rate for a range of $r$ and $K$, and suggests that a small $K$ is generally preferable. This result differs from the practical recommendation in Phillips et al., (2023), where a larger $K$ like 20 is recommended in smaller samples. Because I use a crude union bound to analyze cross-fitting, the rate might be suboptimal regarding $K$. In particular, I might exaggerate the loss in the rate due to a large $K$, so the above analysis might not accurately characterize the impact of $K$ in cross-fitting. A sound theoretical justification on the choice of $K$ is still an open question.
@@ 4 Bias of variance estimators
    For the cross-fit procedure, under Conditions 3, 6 and 8,
    $\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}=\Theta_{+}\left({\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}\right)-\Theta_{+}(K/n).$ (3)
@@
    For the non-cross-fit procedure, under Conditions 3, 9 and $1$,
    $\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}=\Theta_{+}\left({\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}\right)+{\mathrm{O}}(n^{-1/2})-\Theta_{+}(n^{-1}).$ (4)
@@
For the cross-fit procedure, suppose that $K$ is fixed as $n\to\infty$. If ${\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}$ has order $n^{-1}$ (e.g., $\hat{Q}_{k,a}$ is a parametric estimator), then the bias $\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}={\mathrm{O}}(n^{-1})$. If ${\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}={\mathrm{o}}(n^{-1/2})$, as anticipated when $\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}={\mathrm{o}}_{p}(n^{-1/4})$ (a common requirement in the nonparametric inference literature), then the bias $\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}={\mathrm{o}}(n^{-1/2})$ still vanishes at a rate faster than the parametric rate $n^{-1/2}$. If ${\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}$ vanishes slower than $n^{-1}$, then $\sigma_{\dagger,a}^{2}>\sigma_{\#,a}^{2}$ for sufficiently large $n$. That is, for sufficiently large samples, $\hat{\sigma}_{a}^{2}$ is anticipated to overestimate $\sigma_{\#,a}^{2}$, contributing to overcoverage.
@@
Consider the non-crossfit procedure. Suppose that Conditions 3, 9 and 11 hold, and $\tilde{Q}_{a}$ and $Q_{\#,a}$ are projections in the sense that
@@
$\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}={\mathbb{E}}\left[P_{n}\left\{\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)^{2}}\right\}\left\{[Y-\tilde{Q}_{a}(X)]^{2}-[Y-Q_{\#,a}(X)]^{2}\right\}\right]+{\mathbb{E}}[P_{n}(\tilde{Q}_{a}^{2}-Q_{\#,a}^{2})]-\Theta_{+}(n^{-1}),$ (5)
@@
The first term on the right-hand side of (5) is non-positive when $\tilde{Q}_{a}$ fits the data no worse than $Q_{\#,a}$. This holds, for example, if $x\mapsto\pi_{*}(a\mid x)$ is a constant function, $\tilde{Q}_{a}$ minimizes the empirical mean squared error among observations with $A=a$ in the function class $\mathcal{F}$}, and $Q_{\#,a}\in{\mathcal{F}}$. The magnitude of the second term ${\mathbb{E}}[P_{n}(\tilde{Q}_{a}^{2}-Q_{\#,a}^{2})]$ on the right-hand side of (5) can be bounded with bounds for empirical processes, but such analysis might fail to reveal this term’s sign. When $P_{n}\tilde{Q}_{a}^{2}$ is unrelated to the loss for obtaining $\tilde{Q}_{a}$, such bounds might also be overly conservative for its magnitude and ${\mathbb{E}}[P_{n}(\tilde{Q}_{a}^{2}-Q_{\#,a}^{2})]$ might be arguably anticipated to be close to zero. In other cases, $P_{n}\tilde{Q}_{a}^{2}$ often relates to a penalization on the roughness of $\tilde{Q}_{a}$ to avoid overfitting, in which case it might be arguably anticipated that $P_{n}(\tilde{Q}_{a}^{2}-Q_{\#,a}^{2})\leq 0$. Therefore, it might be arguably anticipated that $\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}\leq 0$ if ${\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}$ vanishes slowly and $n$ is large, contributing to decreased CI coverage.
The above analyses also suggest a potential trade-off between efficiency and Wald-CI coverage convergence rate, if the bias of the variance estimator actually contributes substantially to CI coverage, although this trade-off does not show up in (potentially conservative) Theorems 1 and 3. I illustrate with the cross-fit procedure. If we use a highly flexible estimator $\hat{Q}_{k,a}$ of $Q_{*,a}$ in order to achieve more efficiency in the ATE estimator $\hat{\psi}_{a}$, we might anticipate the followings: (i) ${\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}$ vanishes slowly, so (ii) $\sigma_{\dagger,a}^{2}\to\sigma_{\#,a}^{2}$ slowly, and hence (iii) the Wald-CI coverage $\Pr(\sqrt{n}(\hat{\psi}_{a}-\psi_{*,a})\leq z_{\alpha}\hat{\sigma}_{a})$ converges to the nominal coverage $1-\alpha$ slowly.

### v2

@@ 3.2 Non-cross-fit procedure
    Under Part 1 of Condition 14, the left-hand side of (2) is
    ${\mathrm{O}}\left(\Delta_{{\mathrm{VC}},1}\{\log(1/\Delta_{{\mathrm{VC}},1})\}^{1/2}+n^{-1/2}\log(1/\Delta_{{\mathrm{VC}},1})+\left\{(\Delta_{{\mathrm{VC}},1}+n^{-1/2}\right\}\log n+\Delta_{{\mathrm{VC}},2}\right),$
    where $\Delta_{{\mathrm{VC}},1}=({E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2})^{1/3}$ and $\Delta_{{\mathrm{VC}},2}=0$. Additionally under Conditions 11 and 12, if $\tilde{r}_{1}(n)={\mathrm{o}}(\{\log n\}^{-1/\tilde{q}_{1}})$, the rate holds with $\Delta_{{\mathrm{VC}},1}$ and $\Delta_{{\mathrm{VC}},2}$ replaced by $\tilde{r}_{1}(n)(\log n)^{1/\tilde{q}_{1}}$ and $\tilde{r}_{2}(n)(\log n)^{1/\tilde{q}_{2}}$, respectively.
@@
    Under Part 2 of Condition 14, the left-hand side of (2) is
    ${\mathrm{O}}\left(\Delta_{{\mathrm{VC-hull}},1}+\Delta_{{\mathrm{VC-hull}},2}+\left\{\Delta_{{\mathrm{VC-hull}},1}^{(2\nu+1)/(\nu+1)}+n^{-1/2}\right\}\log n\right),$
    where $\Delta_{{\mathrm{VC-hull}},1}=({E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2})^{(\nu+1)/(5\nu+3)}+n^{-(\nu+1)/(6\nu+2)}$ and $\Delta_{{\mathrm{VC-hull}},2}=0$.
    Additionally under Conditions 11 and 12, if ${E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}={\mathrm{o}}(\{\log n\}^{-1/\tilde{q}_{1}})$, this rate holds with $\Delta_{{\mathrm{VC-hull}},1}$ and $\Delta_{{\mathrm{VC-hull}},2}$ replaced by $\{\tilde{r}_{1}(n)\}^{(\nu+1)/(2\nu+1)}(\log n)^{(\nu+1)/\{\tilde{q}_{1}(2\nu+1)\}}$ and $n^{-1/2}\{\tilde{r}_{1}(n)\}^{-2\nu/(2\nu+1)}\allowbreak(\log n)^{-2\nu/\{\tilde{q}_{1}(2\nu+1)\}}+\tilde{r}_{2}(n)(\log n)^{1/\tilde{q}_{2}}$, respectively.
@@ 3.3 Interpretation and comparison of asymptotic rates
For simplicity, particularly when comparing cross-fit and non-cross-fit procedures, I treat $K$ as an absolute constant that does not vary with $n$ for now. Because of the convergence rate of $|\sigma_{\dagger,a}-\sigma_{\#,a}|$ shown in Theorems 2 and 4, the forms in (1) and (2) are somewhat deceiving in the sense that the seemingly first order term $\phi(z_{\alpha})z_{\alpha}(\sigma_{\dagger,a}-\sigma_{\#,a})/\sigma_{\#,a}$ on the left-hand side has a faster convergence rate than the right-hand side, regardless of whether approximate sub-Weibull conditions hold and whether $Q_{\#,a}$ is chosen to yield a faster rate of $|\sigma_{\dagger,a}-\sigma_{\#,a}|$. In other words, the rate of the right-hand side of (1) and (2) is an upper bound on the convergence rate of the Wald-CI coverage to its nominal coverage $1-\alpha$.
@@
For the cross-fit procedure, the ${\mathrm{O}}(\{\log n/n\}^{1/2})$ term in the rate in Theorem 1 is the same as the rate in the Berry-Esseen Theorem except for a log factor. The contribution from estimating the nuisance outcome model can converge slower and strongly depends on whether approximate sub-Weibull conditions 7 and 8 hold: If $\hat{Q}_{k,a}$ is estimated from a (potentially misspecified) parametric model, so that we can take $\hat{r}_{1}(n)=\hat{r}_{2}(n)={\mathrm{O}}(n^{-1/2})$, and approximate sub-Weibull conditions hold, then the contribution from estimating the outcome model is ${\mathrm{O}}(n^{-1/2}\log n)$, same as the Berry-Esseen theorem up to a log factor. However, if $\hat{Q}_{k,a}$ is estimated more flexibly so that $\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}$ has a slower rate, or approximate sub-Weibull conditions do not hold, then the contribution from nuisance estimation can have a slower rate.
@@
For the non-cross-fit procedure, the rate additionally strongly depends on the complexity of the function class $\mathcal{F}$} containing $\tilde{Q}_{a}$, as measured by its covering number. This is because, compared to the cross-fit procedure, Donsker conditions are additionally required to bound empirical process terms, whose rate depends on the function class complexity. Among the empirical process terms in my derivation, the term that impacts the rate most is $(P_{n}-P_{*})\{{\mathcal{T}}_{a}(\tilde{Q}_{a})-{\mathcal{T}}_{a}(Q_{\#,a})\}$ concerning the deviation of $\tilde{\psi}_{a}-\psi_{*,a}$ from the sample mean $(P_{n}-P_{*})D_{a}(Q_{\#,a},\psi_{*,a})$.
@@
By Theorem 3 and Corollary 1, the rates for VC classes (e.g., many parametric models) are similar to the rates for singleton classes, the simplest possible classes, which achieve the lower bound of the right-hand side of (2). These rates are also similar to those for the cross-fit procedure, suggesting that cross-fit and non-cross-fit procedures have similar CI coverage if $\mathcal{F}$} is not rich.
@@
In contrast, the rates for VC-hull classes, which are much richer than VC classes, can be much slower. If the VC index $\nu$ of the associated VC-class is large, namely $\mathcal{F}$} is rich, then the rate would be slow, and the approximate sub-Weibull conditions do not improve the rate as much as in the cross-fit procedure or the non-cross-fit procedure with VC classes. For a rich VC-hull class, namely large $\nu$, the rate contribution from estimating the outcome model is approximately $\tilde{r}_{1}(n)^{1/2}$ up to a log factor, which is slower than the convergence rate of $\tilde{Q}_{a}$. Theorem 3 and Corollary 1 are mainly for illustration rather than for proving accurate rates for specific regression algorithms, because tighter bounds might be derived for a specific algorithm to estimate $Q_{*,a}$ with additional knowledge on this algorithm.
@@ 4 Bias of variance estimators
    For the cross-fit procedure, under Conditions 3, 6 and 9,
    $\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}=\Theta_{+}\left({E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}\right)-\Theta_{+}(K/n).$ (3)
@@
    For the non-cross-fit procedure, under Conditions 3, 10 and 13,
    $\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}=\Theta_{+}\left({E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}\right)+{\mathrm{O}}(n^{-1/2})-\Theta_{+}(n^{-1}).$ (4)
@@
For the cross-fit procedure, suppose that $K$ is fixed as $n\to\infty$. If ${E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}$ has order $n^{-1}$ (e.g., $\hat{Q}_{k,a}$ is a parametric estimator), then the bias $\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}={\mathrm{O}}(n^{-1})$. If ${E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}={\mathrm{o}}(n^{-1/2})$, as anticipated when $\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}={\mathrm{o}}_{p}(n^{-1/4})$ (a common requirement in the nonparametric inference literature), then the bias $\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}={\mathrm{o}}(n^{-1/2})$ still vanishes at a rate faster than the parametric rate $n^{-1/2}$. If ${E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}$ vanishes slower than $n^{-1}$, then $\sigma_{\dagger,a}^{2}>\sigma_{\#,a}^{2}$ for sufficiently large $n$. That is, for sufficiently large samples, $\hat{\sigma}_{a}^{2}$ is anticipated to overestimate $\sigma_{\#,a}^{2}$, contributing to overcoverage.
@@
Consider the non-crossfit procedure. Suppose that Conditions 3, 10 and 13 hold, and $\tilde{Q}_{a}$ and $Q_{\#,a}$ are projections in the sense that
@@
$\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}={E}\left[P_{n}\left\{\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)^{2}}\right\}\left\{[Y-\tilde{Q}_{a}(X)]^{2}-[Y-Q_{\#,a}(X)]^{2}\right\}\right]+{E}[P_{n}(\tilde{Q}_{a}^{2}-Q_{\#,a}^{2})]-\Theta_{+}(n^{-1}),$ (5)
@@
The first term on the right-hand side of (5) is non-positive when $\tilde{Q}_{a}$ fits the data no worse than $Q_{\#,a}$. This holds, for example, if $x\mapsto\pi_{*}(a\mid x)$ is a constant function, $\tilde{Q}_{a}$ minimizes the empirical mean squared error among observations with $A=a$ in the function class $\mathcal{F}$}, and $Q_{\#,a}\in{\mathcal{F}}$. The magnitude of the second term ${E}[P_{n}(\tilde{Q}_{a}^{2}-Q_{\#,a}^{2})]$ on the right-hand side of (5) can be bounded with bounds for empirical processes, but such analysis might fail to reveal this term’s sign. When $P_{n}\tilde{Q}_{a}^{2}$ is unrelated to the loss for obtaining $\tilde{Q}_{a}$, such bounds might also be overly conservative for its magnitude and ${E}[P_{n}(\tilde{Q}_{a}^{2}-Q_{\#,a}^{2})]$ might be arguably anticipated to be close to zero. In other cases, $P_{n}\tilde{Q}_{a}^{2}$ often relates to a penalization on the roughness of $\tilde{Q}_{a}$ to avoid overfitting, in which case it might be arguably anticipated that $P_{n}(\tilde{Q}_{a}^{2}-Q_{\#,a}^{2})\leq 0$. Therefore, it might be arguably anticipated that $\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}\leq 0$ if ${E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}$ vanishes slowly and $n$ is large, contributing to decreased CI coverage.
The above analyses also suggest a potential trade-off between efficiency and Wald-CI coverage convergence rate, if the bias of the variance estimator actually contributes substantially to CI coverage, although this trade-off does not show up in (potentially conservative) Theorems 1 and 3. I illustrate with the cross-fit procedure. If we use a highly flexible estimator $\hat{Q}_{k,a}$ of $Q_{*,a}$ in order to achieve more efficiency in the ATE estimator $\hat{\psi}_{a}$, we might anticipate the followings: (i) ${E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}$ vanishes slowly, so (ii) $\sigma_{\dagger,a}^{2}\to\sigma_{\#,a}^{2}$ slowly, and hence (iii) the Wald-CI coverage ${\mathrm{pr}}(n^{1/2}(\hat{\psi}_{a}-\psi_{*,a})\leq z_{\alpha}\hat{\sigma}_{a})$ converges to the nominal coverage $1-\alpha$ slowly.